OpenStudy (anonymous):

Find a number r such that 0 ≤ r < 9 such that 77 · 139 · 38 · 107 ≡ r ( mod 9).

OpenStudy (anonymous):

do i find the canonical representatives of the numbers in Z9?

Nnesha (nnesha):

hey secret! tag ganeshie8 !!! ;)

OpenStudy (anonymous):

...they are not online

jimthompson5910 (jim_thompson5910):

it might help to reduce every large number as much as possible examples: 77 = 5 (mod 9) since 77/9 = 8 remainder 5 139 = 4 (mod 9) since 139/9 = 15 remainder 4 etc etc

jimthompson5910 (jim_thompson5910):

"do i find the canonical representatives of the numbers in Z9?" Yeah i think so. I'm not familiar with the technical term, but I think that's the same as going from 77 to 5 in mod 9.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i think its 1

jimthompson5910 (jim_thompson5910):

it's not 1

OpenStudy (anonymous):

its 5

jimthompson5910 (jim_thompson5910):

correct